the difference of 23 and 2 times a number is 3 more than 3 times the number. find the number.
i believe the answer is x= 4, is this correct?


Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52785)

.
the difference of 23 and 2 times a number is 3 more than 3 times the number. find the number.
i believe the answer is x= 4, is this correct?
-------------------------------------------------------------------------

23 - 2x = 3x +3,

23 - 3 = 3x + 2x,

5x = 20,

x = 4.

You are right.


Answer by addingup(3677)

23-2x= 3x+3 On both sides, subtract 3 and add 2x
20= 5x Divide both sides by 5
4= x or more elegantly:
x= 4